Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
6x – 4y = 19
24x - 16y = -76

Found 2 solutions by jim_thompson5910, sunsetatdawn:
Answer by jim_thompson5910(35256)

Multiply the first equation by -4 to go from

6x – 4y = 19

to

-24x + 16y = -76


Now add this to the second equation


-24x + 16y = -76
+ 24x - 16y = -76
-----------------------------------
0x + 0y = 0

giving us 0 = 0

Since that last equation is ALWAYS true, this means that there are an infinite number of solutions.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by sunsetatdawn(4)

One slight problem with the previous solution given for this. -76 + -76 is not 0. It is -152. 0 is not equal to -152, meaning this has NO solution. If it were to end with 0=0, then the solution would have infinite solutions, but this is not the case here.